# iMAD: Intelligent Multi-Agent Debate for Efficient and Accurate LLM Inference

### v1

@@ C.1 Ablation Studies: Effectiveness of Structured Self-Critique Prompting
Table 11: Accuracy of beneficial debate decisions using BCE or MSE compared to our FocusCal Loss.
C.2 Ablation Study: Effectiveness of FocusCal Loss
To evaluate the effectiveness of our proposed FocusCal loss, we conduct the evaluation in three aspects: (i) how each loss component of $L_{\text{AF}}$, $L_{\text{CP}}$, and ECE in our FocusCal contributes to the final performance in terms of accuracy and token costs; (ii) how FocusCal outperforms standard losses such as Binary Cross-Entropy (BCE) and Mean Squared Error (MSE) on debate decision correctness; and (iii) what is the impact of the calibration term ECE in FocusCal while replacing only ECE with BCE (li2024rediscoveringbcelossuniform) or MSE (koksoy2006multiresponse) while keeping $L_{\text{AF}}$ and $L_{\text{CP}}$ fixed.
Loss Component Acc (%) # Token
$L_{\text{AF}}$ only 78.8 3,558
$L_{\text{CP}}$ only 78.1 3,379
ECE only 79.1 3,757
$L_{\text{AF}}$ + $L_{\text{CP}}$ 79.7 3,552
$L_{\text{AF}}$ + ECE 79.7 3,561
$L_{\text{CP}}$ + ECE 79.8 3,577
FocusCal ($L_{\text{AF}}$ + $L_{\text{CP}}$ + ECE) 81.3 3,489
FocusCal Loss Terms.
In Table 12, we report accuracy (Acc) and total token cost (# Token) on the VQA-v2 dataset using subsets of the three components in FocusCal (i.e., $L_{\text{AF}}$, $L_{\text{CP}}$, and ECE). We mainly use the VQA-v2 dataset for brevity since the other datasets follow the same performance pattern. When using each single loss term, the behaviors are complementary: $L_{\text{CP}}$ yields the fewest tokens, 3,379 tokens in total, but the lowest accuracy of 78.1% by penalizing overconfident scores raises more examples above the decision threshold, causing the classifier to skip debates, including recoverable ones. ECE attains the best single-term accuracy of up to 79.1% but with the highest token cost of 3,757, as calibration makes scores more faithful to uncertainty and flags more borderline examples for debate, increasing debate frequency even when some debates are unnecessary. $L_{\text{AF}}$ improves accuracy over $L_{\text{CP}}$ by emphasizing hard negatives and lowering scores for wrong-but-confident answers, which triggers helpful debates on those instances.
Two-term combinations consistently improve over single terms. On VQA-v2, using $L_{\text{AF}}~and~\text{ECE}$ and using $L_{\text{CP}}~and~\text{ECE}$ raise accuracy to 79.7% and 79.8% with 3,561 and 3,577 tokens, respectively. This is because ECE term encourages the predicted debate-triggering score to align with empirical correctness, while $L_{\text{AF}}$ penalizes overconfident errors and emphasizes borderline cases, and $L_{\text{CP}}$ penalizes misalignment between the model’s confidence and the semantic uncertainty expressed in the response. In comparison, using $L_{\text{AF}}~and~L_{\text{CP}}$ attains 79.7% with 3,552 tokens. Without ECE, the debate-triggering score is less aligned with empirical correctness, so accuracy is similar but tokens are slightly lower. Using all three terms in FocusCal achieves the highest accuracy (up to 81.3%) with fewer tokens (3,489 tokens), since ECE aligns the debate-triggering score with empirical correctness, $L_{\text{AF}}$ penalizes overconfident errors and emphasizes borderline cases, and $L_{\text{CP}}$ penalizes misalignment between confidence and semantic uncertainty. All three terms together help to better identify recoverable errors and avoid unnecessary debates.
FocusCal vs. BCE/MSE.
We compare our FocusCal loss with two commonly used loss functions to train the binary debate decision classifiers: (i) BCE and (ii) MSE. We measure the percentage of beneficial decisions made by the classifier, defined as correctly skipping debate for correct answers or unrecoverable errors, and correctly triggering debate for recoverable errors. As shown in Table 11, FocusCal achieves the highest rate of beneficial debate decisions across all six datasets, reaching 95.9% on OKVQA and substantially outperforming both BCE (89.3%) and MSE (89.1%). The improvement in beneficial debate decision accuracy stems from FocusCal’s targeted design, which enables the classifier to address two key issues: overconfident incorrect predictions and misalignment between predicted scores and uncertainty signals revealed in the single-agent response. These issues are not effectively handled by BCE or MSE, because they penalize all classification errors equally.
@@ C.2 Ablation Study: Effectiveness of FocusCal Loss
Table 13: Accuracy (Acc) and average token cost (# Token) per question when replacing the ECE term in FocusCal with BCE or MSE while maintaining $L_{\text{AF}}$ and $L_{\text{CP}}$.
Calibration Term of ECE Compared to MSE and BCE.
To further evaluate the design choice of the calibration loss component, we fix $L_{\text{AF}}$ and $L_{\text{CP}}$ and replace the ECE term in FocusCal with BCE or MSE. As shown in Table 13, ECE attains the best or tied accuracy on all datasets and usually uses fewer tokens on average. For example, on GSM8K, ECE improves accuracy over MSE from 82.5% to 84.8% while using about 7% fewer tokens, reducing token usage from 1,099 to 1,025. On VQA-v2, it increases accuracy over BCE from 80.7% to 81.3% and lowers token usage by about 4%, from 3,626 to 3,489. On MMLU, ECE uses a similar number of tokens as MSE (1,010 vs 1,005), but attains higher accuracy, and compared with BCE it improves accuracy while cutting token usage by roughly 22%, from 1,296 to 1,010. Averaged across six datasets, ECE yields 86.4% accuracy with 2,053 tokens, versus 85.85% and 2,297 for BCE, and 85.17% and 2,130 for MSE. These results indicate that ECE calibrates better, whereas BCE and MSE treat errors uniformly and provide weaker calibration, causing borderline cases to either trigger unnecessary debates or skip debates that could recover errors.
C.3 Additional Cross-LLM Evaluation: Qwen 3.0 and GPT-5 nano
@@
Table 14: Performance on GPT-5 nano: Accuracy (Acc) and average token cost (# Token) per question.
@@
Table 15: Performance on Qwen 3.0: Accuracy (Acc) and average token cost (# Token) per question.
We evaluate iMAD on two additional base models, GPT-5 nano (gpt5) and Qwen 3.0 (yang2025qwen3technicalreport), with results in Tables 14 and 15, respectively. All prompt designs, the self-critique stage, memory synchronization, and the debate decision policy are the same as those in the main experimental settings (see Section 5) on Gemini 2.0 Flash. We conducted a one-time sanity check of prompts and thresholds on auxiliary datasets (PubMedQA and GQA) and then froze the debate decision classifier and its threshold. We report both accuracy and the average token cost per question for performance evaluation.
@@
Across both base models, iMAD attains the highest accuracy on every dataset while holding token budgets close to DOWN and far below full-debate methods. On GPT-5 nano with VQA-v2, iMAD reaches 80.1% at 2,258 tokens; GD is 79.0% at 144,917 tokens and MAD is 78.3% at 41,119 tokens. Thus, iMAD matches or exceeds their accuracy while using 98.4% and 94.5% fewer tokens, respectively. On Qwen 3.0 with MMLU, iMAD reaches 91.7% at 1,523 tokens, whereas MAD and GD both achieve 91.3% accuracy but require 149,686 and 309,309 tokens. These correspond to 98.98% and 99.51% lower token cost for iMAD. Full-debate methods incur extremely high token costs because, in every round, they rebroadcast the growing transcript to all agents, expanding the input context even when the example is not recoverable. By contrast, iMAD triggers debate only when the classifier detects hesitation or conflict in the self-critique single-agent output, so extra tokens are spent only on incorrect cases that are likely to be corrected through debate. Moreover, as Tables 14 and 15 show, on reasoning-heavy tasks such as GSM8K, iMAD outperforms single-agent baselines on both models while adding only tens to a few hundred tokens per question. This is because self-critique cues identify cases where triggering debate can fix some errors in intermediate reasoning steps. By contrast, for vision or short factual tasks such as OKVQA, VQA-v2, and MMLU, accuracy gains are smaller, since when the decisive visual cue or fact is missing, triggering additional debate cannot introduce new evidence to flip the incorrect answer.
Compared to DOWN, iMAD achieves higher accuracy with only a small increase in token costs. On GPT-5 nano, iMAD raises MEDQA from 89.0% to 91.6% with 81 additional tokens per question (about 3.9%), GSM8K from 74.8% to 76.5% with 265 more tokens (about 13.2%), and ScienceQA from 87.2% to 89.4% with 180 more tokens (about 8.7%). On Qwen 3.0, iMAD raises MEDQA from 81.5% to 83.8% with 212 more tokens (about 15.8%), GSM8K from 75.5% to 77.6% with 92 more tokens (about 7.2%), and ScienceQA from 87.5% to 90.9% with 51 more tokens (about 4.0%). These gains arise because DOWN relies on the model’s raw confidence, which can be miscalibrated and overconfident, so it often skips debates that are actually needed. iMAD instead utilizes the self-critique responses and uses a FocusCal-trained debate classifier to detect hesitation and conflict, only spend extra tokens on the cases where triggering debate can correct the answer. These accuracy gains arise because DOWN relies on raw model confidence score, which is often miscalibrated and overconfident, leading it to skip debates that are actually needed. In contrast, iMAD first extracts structured features from the self-critique response and then uses a trained debate classifier to decide whether to trigger debate.
With identical prompts and decoding settings, Qwen 3.0 typically generates briefer single-agent responses than GPT-5 nano. This results in lower token budgets in iMAD while preserving similar accuracy gains. Overall, the debate-decision classifier trained once transfers across base models, delivering consistent accuracy improvements with a small token cost while avoiding the heavy cost of full debate.
@@ Appendix D Discussion
While iMAD demonstrates strong improvements in both accuracy and token efficiency, here we discuss its design assumptions and potential extension.
@@
One key requirement for iMAD is constructing a training dataset with correctness labels for single-agent answers, which we create by prompting a single agent on two representative datasets (i.e., one QA and one VQA) and manually labeling each response as correct or incorrect. These binary correctness labels serve as supervision for training the debate decision classifier: each example is paired with features extracted from the self-critique response, and the classifier learns to predict whether a debate would be beneficial. Although this labeling process is lightweight and limited to only two datasets, our careful dataset selection exposes the classifier to diverse reasoning behaviors. This design allows the classifier to generalize well and perform effectively on six held-out datasets without further tuning. However, because the classifier is trained offline and remains fixed during deployment, it may not effectively adapt to changes in model behavior. Future work could explore online adaptation through feedback-driven updates or weak supervision to further improve generalization.
@@
Another consideration lies in the reliance on structured self-critique prompting, which iMAD uses to extract interpretable features from the model’s output to train the downstream debate classifier to capture general model behavior. However, this design assumes that the underlying LLM agent can effectively articulate its reasoning and express uncertainty. In domains where the model’s responses are incoherent or fail to convey uncertainty clearly, the extracted features may become less informative. In such cases, additional prompt refinement may be necessary to preserve debate decision quality.
@@
An interesting extension is to move the debate–triggering decision inside the generation process rather than waiting for the full self-critique to complete. In our current deployment, the LLM agent is accessed only through a black-box API (e.g., Gemini or GPT), so iMAD can decide whether to trigger debate only after the complete self-critique is returned. With open-source models or streaming APIs, we can instrument token-by-token generation, monitor the emerging rationale for early hesitation cues, and trigger MAD before the self-critique finishes, potentially reducing both latency and token costs while maintaining accuracy. A further direction is to combine these in-generation signals with internal model states (such as logit trajectories or entropy over the candidate answers) and learn an adaptive policy (e.g., reinforcement learning), that allocates debate budget dynamically over the course of generation.

### v2

@@ C.2 Ablation Study on FocusCal Loss
Table 12: Percentage of beneficial debate decisions using BCE or MSE compared to our FocusCal Loss.
@@
Table 13: Accuracy (Acc) and average token cost (# Token) per question when replacing the ECE term in FocusCal with BCE or MSE while maintaining $L_{\text{AF}}$ and $L_{\text{CP}}$. Bold values indicate the best result in each row, and underlined values indicate the second best.
FocusCal Loss Terms.
In Table 11, we report accuracy (Acc) and total token cost (# Token) on the VQA-v2 dataset when using different subsets of the three components in FocusCal, i.e., $L_{\text{AF}}$, $L_{\text{CP}}$, and ECE. We report the results on the VQA-v2 dataset for brevity, since the other five datasets exhibit a similar performance pattern in both accuracy and token costs (e.g., full FocusCal consistently achieves higher accuracy than any two-term combination, which in turn outperforms any single-term loss). When we use only a single loss term, each component shows a distinct behavior. For example, using $L_{\text{CP}}$ alone yields the lowest token cost (3,379 tokens in total) but also the lowest accuracy of 78.1%. Since $L_{\text{CP}}$ penalizes overconfident scores, it tends to push many predicted debate-triggering scores below the decision threshold. As a result, the classifier skips debate on more questions, including some that would have benefited from MAD, which reduces accuracy. Using ECE alone attains the best single-term accuracy of 79.1% but with the highest token cost of 3,757 tokens. The ECE term adjusts the debate-triggering scores so that instances with scores close to the decision threshold (i.e., cases where the classifier is not clearly confident about skipping or triggering debate) are more likely to trigger MAD. This causes the classifier to trigger MAD on more questions, which improves accuracy but also increases token usage, even though some of these extra debates are unnecessary. Using $L_{\text{AF}}$ alone improves accuracy over $L_{\text{CP}}$ by emphasizing hard error cases (i.e., lowering the scores for wrong-but-confident answers). This behavior encourages the classifier to trigger debates precisely on those difficult cases, which helps correct more errors.
Combining two loss terms consistently improves performance over single-term configurations since each loss term targets a different failure mode of the debate decision classifier. On VQA-v2, the combination of $L_{\text{AF}}$ and ECE and the combination of $L_{\text{CP}}$ and ECE both raise accuracy to 79.7% and 79.8%, with token costs of 3,561 and 3,577, respectively. In these settings, ECE encourages the debate-triggering scores to match empirical correctness, while $L_{\text{AF}}$ focuses on overconfident errors and on examples near the decision threshold (where the classifier is unsure whether to skip or trigger debate), and $L_{\text{CP}}$ penalizes mismatches between the predicted score and the semantic uncertainty expressed in the self-critique response.
For comparison, using $L_{\text{AF}}$ and $L_{\text{CP}}$ without ECE attains 79.7% accuracy with 3,552 tokens, which is lower accuracy and slightly higher token cost than FocusCal. Removing the ECE term makes the debate-triggering scores less calibrated, so the classifier skips some recoverable errors (reducing accuracy) and also sends additional already correct or unrecoverable questions to debate (increasing token usage). Removing the ECE term makes the debate-triggering scores less calibrated, causing the classifier to skip some recoverable errors and thus reduce accuracy, and to trigger MAD on more already correct or unrecoverable instances and thus increase token costs.
Using all three terms together in FocusCal achieves the best overall performance on VQA-v2, reaching the highest accuracy of 81.3% while reducing token usage to 3,489 tokens. In FocusCal, ECE calibrates the debate-triggering scores toward empirical correctness, $L_{\text{AF}}$ downweights wrong-but-confident predictions and highlights difficult cases, and $L_{\text{CP}}$ enforces consistency between predicted scores and semantic uncertainty cues in the self-critique. All three terms in FocusCal help the classifier identify more recoverable errors while avoiding unnecessary debates, which leads to higher accuracy and lower token usage.
FocusCal vs. BCE and MSE.
We compare FocusCal with two commonly used losses for training binary classifiers: BCE and MSE. Here, we focus on the quality of debate decisions rather than raw accuracy. Specifically, we measure the percentage of beneficial decisions made by the classifier, defined as the percentage of examples where the classifier (i) correctly skips debate for questions that are already correct or not recoverable by MAD, or (ii) correctly triggers debate for questions where MAD can improve the answer.
As shown in Table 12, FocusCal achieves the highest rate of beneficial decisions across all datasets. For example, on OKVQA, FocusCal achieves 95.9% beneficial decisions, while BCE and MSE achieve 89.3% and 89.1%, respectively. The improvement comes from the targeted design of FocusCal. By combining $L_{\text{AF}}$, $L_{\text{CP}}$, and ECE, FocusCal directly targets two key issues that matter for debate decisions: overconfident incorrect predictions and mismatches between the predicted debate-triggering scores and the uncertainty signals in the self-critique single-agent response. In contrast, BCE and MSE provide weaker guidance on when a debate is truly warranted, as they treat all classification errors uniformly and do not explicitly address these issues.
@@
Table 14: Performance on GPT-5 nano: Accuracy (Acc) and average token cost (# Token) per question. Bold values indicate the best accuracy in each row, and underlined values indicate the second best.
@@
Table 15: Performance on Qwen 3.0: Accuracy (Acc) and average token cost (# Token) per question. Bold values indicate the best accuracy in each row, and underlined values indicate the second best.
Calibration Term of ECE Compared to MSE and BCE.
To further evaluate the design choice of the calibration component ECE in FocusCal, we fix $L_{\text{AF}}$ and $L_{\text{CP}}$ and only replace the ECE term with BCE or MSE. As shown in Table 13, the model trained with ECE attains the best or tied accuracy on all datasets and usually uses fewer tokens. For example, on GSM8K, using ECE instead of MSE improves accuracy from 82.5% to 84.8% while reducing token usage by about 7% (from 1,099 to 1,025 tokens per question). A similar pattern appears on VQA-v2, where ECE increases accuracy over BCE from 80.7% to 81.3% and lowers token usage by about 4% (from 3,626 to 3,489 tokens). Averaged across all six datasets, training with ECE yields 86.4% accuracy with 2,053 tokens per question, whereas BCE yields 85.9% accuracy with 2,297 tokens, and MSE yields 85.2% accuracy with 2,130 tokens.
These results show that using ECE as the calibration component works better than using BCE or MSE in this setting. ECE groups instances into bins according to their predicted debate-triggering scores. For each bin, it compares the average predicted score with the actual fraction of instances in that bin where triggering debate is beneficial. When these two numbers differ a lot (e.g., when the bin has high scores but MAD rarely helps, or low scores but MAD often helps), the ECE term becomes large, pushing the model to adjust its scores in that bin. In contrast, BCE and MSE update each instance independently and do not explicitly fix these group-level mismatches. As a result, they are more likely to assign scores that are too high to many non-beneficial cases or too low to many beneficial cases, which makes the fixed decision threshold less reliable.
C.3 Cross-LLM Evaluation
To evaluate the cross-model performance of iMAD, we conduct an additional cross-LLM study from three perspectives: (i) comparison of the token costs and accuracy of iMAD against all baselines when switching the base model to GPT-5 nano (gpt5) and Qwen 3.0 (yang2025qwen3technicalreport); and (ii) analysis of how differences between base models influence the accuracy and token cost of iMAD.
Experimental Setup.
We evaluate iMAD on the six benchmarks for both GPT-5 nano and Qwen 3.0, with detailed results reported in Table 14 for GPT-5 nano and Table 15 for Qwen 3.0. All prompt templates, including the self-critique stage and memory synchronization, are kept identical to the main experimental setting with Gemini 2.0 Flash (Section 5). For each base model, we first regenerate self-critique responses on the same two auxiliary datasets used in the main experiments (PubMedQA for QA and GQA for VQA), and retrain the debate decision classifier with FocusCal using these responses and their labels. We then select a decision threshold on these auxiliary datasets and keep both the classifier and the threshold fixed for that base model during evaluation. After this training step, we apply iMAD to all six evaluation datasets without any further tuning. For each method and each dataset, we report accuracy and the average total number of tokens per question (sum of input and output tokens) to evaluate performance.
Token Cost and Accuracy across Base Models.
@@ C.3 Cross-LLM Evaluation
Across both GPT-5 nano and Qwen 3.0, iMAD attains the highest accuracy on every dataset, while keeping token usage close to DOWN and far below full-debate methods. On GPT-5 nano with VQA-v2, iMAD achieves 80.1% accuracy with 2,258 tokens per question, whereas GD reaches 79.0% accuracy with 144,917 tokens, and MAD reaches 78.3% accuracy with 41,119 tokens. Thus, iMAD matches or exceeds the accuracy of full-debate methods while using 98.4% and 94.5% fewer tokens than GD and MAD, respectively. On Qwen 3.0 with MMLU, iMAD attains 91.7% accuracy with 1,523 tokens per question, while MAD and GD both reach 91.3% accuracy but require 149,686 and 309,309 tokens. In this case, iMAD reduces token cost by 98.98% relative to MAD and by 99.51% relative to GD. Full-debate methods incur such high token costs because they trigger MAD on every instance, regardless of whether the instance is easy, already correctly answered, or unrecoverable. In every round, they append new messages to the shared transcript and resend the entire cumulative transcript to all agents. This repeated rebroadcast quickly increases the number of input token costs, leading to very large token budgets even on instances where debate does not change the final answer. In contrast, iMAD triggers debate only when the debate decision classifier detects hesitation or conflict in the self-critique single-agent response. As a result, extra tokens are spent mainly on the questions where triggering MAD is likely to correct an error.
Compared with DOWN, iMAD consistently achieves higher accuracy with only a slight increase in token usage on both base models. On GPT-5 nano, for example, iMAD improves MEDQA accuracy from 89.0% to 91.6% while using 81 more tokens per question (about 3.9% more tokens). On Qwen 3.0, iMAD improves MMLU accuracy from 88.9% to 91.7% with 53 additional tokens per question (about 3.6% more tokens). These results are consistent with the main experiments on the Gemini 2.0 Flash model and show that the advantage of iMAD over DOWN persists across base models. These gains arise from the different ways the two methods decide when to trigger debate. DOWN relies solely on the model-generated confidence score in the single-agent answer and compares this score against a dataset-specific threshold. As confidence scores from LLMs are often overconfident even when the answer is incorrect, DOWN frequently skips debate on instances where additional debate would be beneficial. In contrast, iMAD uses a dedicated debate decision classifier trained with the FocusCal loss to decide whether to trigger MAD. This classifier learns to trigger MAD when the self-critique shows hesitation or conflicting reasoning, and to skip debate when the self-critique strongly supports the initial correct answer or indicates that the error is unlikely to be fixed by debate. As a result, iMAD spends additional tokens mainly on instances where triggering MAD is likely to correct the error, which explains why it achieves higher accuracy than DOWN with only a slightly higher token usage.
Furthermore, on GSM8K, a reasoning-heavy math word-problem dataset that requires multi-step numerical reasoning, iMAD improves over single-agent baselines on both GPT-5 nano and Qwen 3.0 while adding only a small number of extra tokens per question. In these settings, many errors arise from intermediate reasoning steps. The self-critique stage therefore elicits additional reasoning that reveals uncertainty and alternative reasoning paths, which gives the debate decision classifier clearer signals about when triggering MAD is likely to correct the error. For VQA or short factual tasks such as OKVQA, VQA-v2, and MMLU, the accuracy gains are smaller because these tasks often hinge on a single decisive visual cue or fact. When that key cue is hard to obtain, triggering MAD adds little new information and has limited ability to correct the error.
Effect of Base Model Differences.
Under the same prompts and overall experimental setup, we observe that iMAD uses fewer tokens per question on Qwen 3.0 than on GPT-5 nano (see Table 15 for Qwen 3.0 and Table 14 for GPT-5 nano). This difference reflects the combined effect of each model’s tokenizer and how long the models typically respond under the same prompts. As a result, iMAD achieves similar accuracy on both models while Qwen 3.0 uses fewer tokens per question than GPT-5 nano. This suggests that iMAD benefits more from base models whose responses are compact but still informative. In this case, fewer tokens are spent per question, while the debate decision classifier still receives enough useful signals, such as uncertainty or conflicting reasoning, to decide whether triggering MAD is beneficial.
Overall, these results show that, for each base model, the debate decision classifier trained only once on the auxiliary datasets (e.g., PubMedQA and GQA) generalizes well to the six evaluation datasets. Across both GPT-5 nano and Qwen 3.0, iMAD consistently improves accuracy over single-agent baselines and over DOWN, while avoiding the heavy token cost of full-debate methods.
@@ Appendix D Discussion
While iMAD demonstrates its strong improvements in both token efficiency and accuracy, we discuss its design assumptions (i.e., limited labeled data, prompt engineering, and label design choices for the debate classifier) and potential extensions for future work.
@@
One key requirement for iMAD is constructing a training dataset with correctness labels for single-agent answers. We construct this dataset by running the self-critique single-agent pipeline on two auxiliary datasets, PubMedQA (a QA dataset) and GQA (a VQA dataset), and manually checking whether the chosen option matches the ground-truth answer. We label the single-agent answer as correct when its chosen option matches the ground-truth answer, and as incorrect otherwise. These binary correctness labels serve as supervision for training the debate decision classifier: each example is paired with features extracted from the self-critique response, and the classifier learns to predict whether a debate would be beneficial. This labeling process is lightweight and uses only two datasets, and our careful dataset selection exposes the classifier to diverse reasoning behaviors. This design allows the classifier to generalize well and perform effectively on six held-out datasets without further tuning. However, since the classifier is trained offline and remains fixed during deployment, it may not effectively adapt to changes in model behaviors. Future work could explore online adaptation through feedback-driven updates or weak supervision to further improve generalization.
Although the labeling relies on only two datasets, they cover diverse question types and reasoning patterns, which helps the classifier generalize to six held-out evaluation datasets without additional tuning. A limitation is that the classifier is trained offline and kept fixed at deployment, so it cannot adapt when the underlying model behaviors drift over time or when the deployment domain changes substantially. Future work could explore online adaptation, for example, through feedback-driven updates or weak supervision, to further improve generalization.
@@
Another consideration lies in the reliance on structured self-critique prompting, which iMAD uses to extract interpretable features from the model’s output to train the downstream debate decision classifier to capture general model behaviors. However, this design assumes that the underlying LLM agent can effectively articulate its reasoning and express uncertainty. In domains where the model’s responses are incoherent or fail to convey uncertainty clearly, the extracted features may become less informative. In such cases, additional prompt refinement may be necessary to preserve debate decision quality.
Label Design Choices for the Debate Classifier.
For the debate decision classifier, we explored two label design choices. The first, which we adopt in iMAD, is a simple correctness-based labeling strategy. We construct a training dataset by running the self-critique single-agent pipeline on two auxiliary datasets (PubMedQA for QA and GQA for VQA) and comparing the single-agent answer with the ground-truth label for each instance, as described in Section 4.1. The second strategy is a more direct debate-benefit labeling scheme that labels each instance according to whether MAD actually improves the single-agent answer. In this debate-benefit scheme, we run both the single-agent and MAD pipelines on the training instances and assign a binary label $y$ to each instance. We set $y=1$ if the single-agent answer is wrong and the MAD answer is correct, so debate should be triggered. In all other cases, including when both answers are correct, both are wrong, or the single-agent answer is correct, but the MAD answer is wrong, we set $y=0$, since debate is unnecessary or harmful.
In practice, this debate-benefit labeling scheme performed poorly for training the debate decision classifier, for two main reasons. First, the positive cases with $y=1$, where MAD actually fixes a wrong single-agent answer, are much fewer than the negative cases (about 6-9% of all instances), so the training dataset is highly imbalanced. The classifier therefore tends to learn a trivial policy that rarely triggers debate. Second, the labels produced by this scheme do not align well with what the classifier actually observes. In the debate-benefit scheme, each label records whether MAD corrects the single-agent error in that particular run. However, MAD’s success depends not only on the single-agent output, but also on the detailed debate trajectory and sampling randomness during generation. For the same question and an almost identical self-critique single-agent response, one MAD run may fix the answer while another may fail. The debate decision classifier, by design, only uses features derived from the single-agent self-critique and never has access to the MAD trajectory when it makes its prediction. Consequently, two training instances with nearly identical single-agent features can receive opposite labels, depending on how MAD happens to behave in that run. From the classifier’s perspective, the same input pattern is sometimes labeled “trigger” and sometimes labeled “skip”. As a result, the labels are noisy and inconsistent with the input features, and the classifier cannot learn a clear rule for when triggering MAD is beneficial.
Overall, these observations lead us to adopt the correctness-based labeling strategy in our final design: it is both cheaper to obtain and more robust for training the debate decision classifier. In contrast, constructing debate-benefit labels requires running the full MAD pipeline for every training example, which is far more expensive in token cost than the correctness-based labeling pipeline. In iMAD, the training labels depend only on whether the single-agent answer is correct or not. Hence, they are robust and do not involve any randomness from the debate process during labeling. Randomness in MAD appears only at inference time and does not affect how the classifier is trained. Training with binary correctness labels, together with the FocusCal loss, encourages the classifier to trigger MAD mainly on instances where the self-critique single-agent response reveals high uncertainty or self-contradictory. These cases strongly indicate single-agent errors that MAD can often correct. This does not mean that every incorrect single-agent answer will trigger MAD. Instead, the classifier focuses on those errors that can be detected from the self-critique response and for which debate is more likely to be beneficial.
@@
An interesting extension is to make the debate-triggering decision during output generation rather than waiting for the entire self-critique response to be generated. In our current deployment, the LLM agent is accessed only through a black-box API (e.g., Gemini or GPT), so iMAD can decide whether to trigger debate only after the complete self-critique response is generated. With open-source models or streaming APIs, it would be possible to monitor token-by-token generation and detect early signs of either hesitation or strong confidence in the emerging rationale. Such signals could be used to trigger MAD early when needed or skip debate when the answer appears well justified, potentially reducing both inference time and token costs while maintaining accuracy. A further direction is to combine such in-generation signals with internal model states (e.g., logit trajectories or entropy over the candidate answers) and learn an adaptive policy (e.g., via reinforcement learning) that allocates debate budget dynamically over the course of generation.